verify that the indicated function is a solution of the given differential equation. Assume an appropriate interval I of denition of each solution. 25. x dy dx 2 3xy 5 1; y 5 e3x # x 1 e23t t d


1
Expert's answer
2021-10-27T13:30:18-0400
"x\dfrac{dy}{dx}-3xy=1"

"y=e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt"

"\dfrac{dy}{dx}=\dfrac{d}{dx}\bigg(e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt\bigg)"

"=3e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt+e^{3x}\dfrac{e^{-3x}}{x}"

"=3e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt+\dfrac{1}{x}"

Substitute


"x\dfrac{dy}{dx}-3xy=1"

"x\bigg(3e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt+\dfrac{1}{x}\bigg)"

"-3x\bigg(e^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt\bigg)=1"

"3xe^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt+1-3xe^{3x}\displaystyle\int_{1}^{x}\dfrac{e^{-3t}}{t}dt=1"

"1=1, True"

Therefore the indicated function is a solution of the given differential equation.